OpenStudy (anonymous):

urgently needed...plseeeeeeeeeee prove n^3-n is always divisible by 8

OpenStudy (anonymous):

plse solve it without by mathematical induction

terenzreignz (terenzreignz):

n = 2 is a contradiction?

OpenStudy (mertsj):

It's not true.

OpenStudy (mertsj):

2^3-2 =6 4^3-4=60 6^3-6=210

terenzreignz (terenzreignz):

In fact, false for all even numbers which aren't divisible by 8.

OpenStudy (anonymous):

i know , i also put all all the values like u have put, but how to solve step by step plsee eplain

terenzreignz (terenzreignz):

It's simply not true. One counterexample is enough (n = 2)

OpenStudy (anonymous):

okay.. @terenzreignz and @Mertsj thank u

OpenStudy (mertsj):

Are there restrictions on n that you did not post?

OpenStudy (anonymous):

no, there was no restrictions, only wriiten, for a positve integer n, prove it